OpenStudy (anonymous):

Bella, Lydia, Camila, and Trinity together stacked a total of 60 pizzas in a pan. Bella stacked 0.25 of the pizzas, Lydia stacked 12 of the 60 pizzas, Camila stacked 35% of the pizzas, and Trinity stacked the remaining pizzas. Who stacked the most pizzas?

OpenStudy (welshfella):

first convert the percentages and fractions into number of pizzas for example Bella stacked 0.25 * 60 pizzas

OpenStudy (welshfella):

what does 0.25 * 60 work out to?

OpenStudy (anonymous):

15 @welshfella

OpenStudy (welshfella):

right and for Camila:- 35% of 60 = 35 * 60 2100 ----- = ---- 100 100

OpenStudy (welshfella):

so that is how many pizzas?

OpenStudy (anonymous):

21

OpenStudy (welshfella):

OK so we have 12 , 15 and 21 pizzas so Trinity stacked 60 - 12 - 15 - 21 right?

OpenStudy (anonymous):

right

OpenStudy (welshfella):

can you work that out then we can see which person stacked the most

OpenStudy (welshfella):

lets add up 12 + 15 + 21 that comes to 48 so trinity stacked 60 - 48 = 12

OpenStudy (welshfella):

So who stacked the most?

OpenStudy (anonymous):

lydia